OpenStudy (anonymous):

Use the 1-1 property of the exponential function to solve the given equation. a. 3^-x=27 This is what I worked out. log3(3^-x)=log3(27) -x=log(27)/log3 -x=3 Am I on the right track or no?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

-x = 3, thus x = -3

OpenStudy (anonymous):

Thank you!!

OpenStudy (jdoe0001):

\(\bf 3^{-x}=27\implies \cfrac{1}{3^x}=27\implies \cfrac{1}{27}=3^x\quad 27=3^3\quad thus\\ \quad \\ \cfrac{1}{3^3}=3^x\implies 3^{-3}=3^x\)